# MMReview: A Multidisciplinary and Multimodal Benchmark for LLM-Based Peer Review Automation

### v2

@@ Abstract
With the rapid growth of academic publications, peer review has become an essential yet time-consuming responsibility within the research community. Large Language Models (LLMs) have increasingly been adopted to assist in the generation of review comments; however, current LLM-based review tasks lack a unified evaluation benchmark to rigorously assess the models’ ability to produce comprehensive, accurate, and human-aligned assessments—particularly in scenarios involving multimodal content such as figures and tables. To address this gap, we propose MMReview, a comprehensive benchmark that spans multiple disciplines and modalities. MMReview includes multimodal content and expert-written review comments for 240 papers across 17 research domains within four major academic disciplines: Artificial Intelligence, Natural Sciences, Engineering Sciences, and Social Sciences. We design a total of 13 tasks grouped into four core categories, aimed at evaluating the performance of LLMs and Multimodal LLMs (MLLMs) in step-wise review generation, outcome formulation, alignment with human preferences, and robustness to adversarial input manipulation. Extensive experiments conducted on 16 open-source models and 5 advanced closed-source models demonstrate the thoroughness of the benchmark. We envision MMReview as a critical step toward establishing a standardized foundation for the development of automated peer review systems.
Introduction
Peer review is essential to scholarly publishing, ensuring research quality and enhancing academic writing. However, the growing volume of submissions has strained the traditional review process, leading to inefficiencies and limited reviewer availability (Kim, Lee, and Lee 2025), which restricts feedback and delays review outcomes. Advances in LLMs have made automated peer review increasingly viable, as these models show strong reasoning abilities and can offer constructive feedback on academic manuscripts (Liu and Shah 2023; Zhao et al. 2024; Zhuang et al. 2025), partially alleviating reviewer burden. Yet, current evaluations of LLM-generated reviews focus mainly on final outputs, lacking in-depth analysis of the reasoning processes behind model judgments. Additionally, most studies concentrate on AI papers with publicly available text, overlooking the multimodal nature of academic papers—such as figures and tables—and the evaluation of LLMs in reviewing research across broader scientific domains.
@@ Introduction
  • We conduct extensive experiments on 16 open-source and 5 closed-source models using the MMReview benchmark, offering some key findings of LLM-based automated reviewing. Our findings offer in-depth analysis and valuable guidance for the future development of LLM-assisted peer review systems.
Related Works
LLMs for Paper Review
LLMs have shown strong potential in analyzing complex scholarly texts (Liu and Shah 2023; Zhao et al. 2024; Zhuang et al. 2025). Initial studies indicate that LLM-generated review comments partially overlap with those of human reviewers, suggesting their potential contribution to peer review (Robertson 2023; Liang et al. 2023). However, further research reveals that even advanced models like GPT-4o often fail to meet human expectations in review quality (Zhou, Chen, and Yu 2024). To improve alignment with peer review standards, researchers have built datasets from public review platforms and fine-tuned LLMs (Kang et al. 2018; Yuan, Liu, and Neubig 2021; Shen et al. 2022; Dycke, Kuznetsov, and Gurevych 2023; Gao, Brantley, and Joachims 2024). Other methods involve multi-turn dialogue (Tan et al. 2024) or multi-agent prompting (D’Arcy et al. 2024) to generate comprehensive feedback. Nonetheless, these efforts focus solely on textual reviews, neglecting the vital role of multimodal content—such as figures and tables—and lack rigorous analysis of the reasoning processes behind LLM-generated critiques.
Evaluation for LLM-based Peer Review
Prior studies (Shen et al. 2022; Yu et al. 2024; Gao, Brantley, and Joachims 2024; Tan et al. 2024; Gao et al. 2025) have predominantly evaluated the quality of LLM-generated peer review comments by measuring their correlation or similarity with human-written reviews using automated metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), BERTScore (Zhang et al. 2020), and METEOR (Banerjee and Lavie 2005). In addition, several studies (Robertson 2023; Zhou, Chen, and Yu 2024; Gao et al. 2025) have adopted the LLM-as-a-judge paradigm, leveraging cutting-edge language models to assess the quality of review comments produced by other LLMs. Given the absence of an established gold standard for this evaluation task, recent research (Xu et al. 2024) has introduced the Generative Estimator for Mutual Information (GEM) to quantify the degree of semantic overlap between LLM-generated and human-authored reviews. Nevertheless, existing evaluation methodologies are not grounded in a unified benchmark or task framework, and they fall short of providing a comprehensive analysis of the underlying reasoning processes involved in LLM-generated peer review.
@@
Overall Pipeline of MMReview Benchmark
@@
Data Processing
@@
During the data processing stage, we first filtered the collected seed dataset $D$ to obtain high-quality papers while maintaining a relatively balanced distribution across disciplines and ensuring a comparable number of accepted and rejected papers. To achieve this, we designed a dual-path joint data selection mechanism that simultaneously filters for paper quality and enforces distributional balance.
@@
Specifically, we utilized Qwen3-32B [^1] Qwen3-32B ranks among the top 10 on the OpenCompass leaderboard, offering a favorable trade-off between performance and model size, making it efficient for large-scale sample filtering (Yang et al. 2025) to generate summaries of each paper under two distinct input conditions: one using only the abstract and the other using the full text of the manuscript. A greater divergence between the two generated summaries is interpreted as evidence that the full text provides substantially more information, thus indicating higher paper quality. Samples with a significant information gain from the full text were retained in our test benchmark. To further ensure that the benchmark maintains a balanced distribution across academic disciplines and an approximately equal ratio of accepted to rejected papers, we supplemented the benchmark by manually incorporating top-ranked papers from specific domains in $D$ based on the quality rankings. This human-filtering procedure guarantees that the composition of the benchmark dataset aligns with our desired distributional properties.
@@
During the data filtering phase, we obtained a total of 240 paper samples spanning 17 research domains across 4 major disciplines to construct our evaluation benchmark, denoted as $D_{b}$. The statistical details of $D_{b}$, including the number of samples per domain and the distribution of accepted versus rejected papers, are presented in Figure 1. For each of these 240 samples, we utilized PDF parsing tools to extract textual content, figures, and tables from the manuscript files, and converted each page of the PDF into corresponding images. As a result, we constructed three distinct modalities of input data: text-only, multimodal (text combined with extracted visual elements), and PDF-as-image.
Task Construction
@@
During task construction, we developed 13 tasks grouped into four thematic categories—step-based, outcome-based, preference-based, and attack-based—reflecting the peer review workflow and challenges LLMs may face. Task distribution is shown in Figure 1. Prompts were designed based on reviewer guidelines from major academic conferences. For each task, we used regular expressions or GPT-4o to extract reference answers from human reviews in $D_{b}$, forming the question–answer dataset $D_{QA}$. For the Fake Strength Evaluation and Fake Weakness Evaluation tasks, GPT-4o and custom prompts generated antonymic rewrites, which were manually verified for semantic accuracy. This process finalized the MMReview benchmark. Figure 2 illustrates its multidisciplinary scope, multimodal inputs, and diverse task design. The detailed prompts of task generation and each task can be found in the supplementary materials.
Step-based Tasks
@@
Summarizing a paper is the initial step in peer review and a key test of a model’s ability to extract essential content. Inaccurate summaries may impair subsequent review generation. To address this, the Summary task assesses a model’s ability to distill key information from a full manuscript into an accurate, concise summary. The model generates a brief summary in its own words, avoiding abstract copying and subjective judgment. These outputs are compared to human-written summaries and evaluated for semantic similarity and information coverage, measuring the model’s holistic comprehension and representation of academic content.
@@
The Soundness Scoring and Presentation Scoring tasks evaluate LLMs’ ability to quantitatively assess manuscript quality, focusing on technical soundness and writing presentation. In Soundness Scoring, the model rates the reliability of technical claims, experimental rigor, and evidential support, emphasizing empirical and methodological validity. Presentation Scoring assesses linguistic clarity and logical organization, reflecting writing quality and information structure. Both tasks require integer scores from 1 to 4, denoting “poor” to “excellent.” Model scores are compared to human ratings to assess judgment consistency.
Figure 2: An overview of MMReview benchmark.
Outcome-based Tasks
@@
The Direct Decision and CoT (Chain-of-Thought) Decision tasks evaluate LLMs’ ability to autonomously assess academic paper quality, reflecting two reviewing paradigms: streamlined judgment and step-by-step reasoning (Wei et al. 2022). These tasks vary in input format and cognitive complexity, enabling controlled comparison of model performance under different reasoning demands. In the Direct Decision task, the model produces an overall score without guidance, simulating a reviewer’s holistic judgment from a single read. In contrast, the CoT Decision task guides the model through a structured reasoning process: summarizing the paper, analyzing strengths and weaknesses across Quality, Clarity, Significance, and Originality, assigning Soundness and Presentation scores, and synthesizing an overall score. This mirrors a reviewer’s iterative, analytical evaluation. The tasks test reasoning ability, analytical structure, and scoring traceability. Model outputs are compared to human scores to assess consistency and rationality. While the Direct task emphasizes efficiency, the CoT task offers interpretability, together forming a complementary framework for evaluating LLMs’ scoring competence.
@@
Beyond generating individual reviews and scores, a key aspect of academic peer review is the Area Chair’s (AC) synthesis of reviewer feedback to make a final decision. To emulate this, the Meta Decision task requires the model to issue a binary judgment—Accept or Reject—based on multiple human reviews, testing its ability to aggregate information and balance differing views. The prompt provides structured guidance and evaluation criteria, prompting step-by-step reasoning. The model is instructed to assess the quality and consistency of reviews rather than merely averaging scores. This task mirrors the real role of an AC and rigorously evaluates the model’s capacity for synthesis, decision rationality, and transparency. Model outputs are compared to human-written meta-reviews to assess reliability and scientific judgment in high-level peer review.
Preference-based Task: Pairwise Rank (PR)
Prior work has shown that pairwise comparison effectively evaluates the alignment between LLM-generated preferences and human judgments (Liu et al. 2024b). Since academic conference acceptance tiers—oral, spotlight, poster, reject—reflect human preference rankings, the Pairwise Rank task is designed to test whether LLMs, as reviewers, display preference patterns consistent with human evaluators. This task assesses the model’s relative judgment ability by presenting pairs of papers from different acceptance tiers: oral (top 5%), spotlight (top 25%), poster, and reject. The model compares and ranks the papers, simulating real-world peer review selection. Alignment is measured by comparing model preferences with actual acceptance categories to determine ranking accuracy. To reduce positional bias (Shi et al. 2025; Thakur et al. 2025), each comparison is repeated with reversed input order.
Attack-based Tasks
This task category assesses models’ robustness and discriminative ability in peer review by introducing adversarial inputs. Fabricated strengths or weaknesses—such as inverted pros and cons—are used to test the model’s capacity to detect factual inaccuracies. Misleading prompts further aim to divert the model from its original instructions. Model outputs are compared to human judgments or the model’s own non-adversarial responses, focusing on its ability to preserve evaluative independence and logical consistency under input perturbations.
@@
LLMs have been widely observed to exhibit a tendency toward “sycophancy”—an inclination to conform to human judgments (Liu et al. 2024a; Kumaran et al. 2025)—which may impair their ability to identify flaws and generate critical peer review comments. Furthermore, the hallucination problem inherent in LLMs can result in inaccurate evaluations of a paper’s strengths and weaknesses. To investigate these issues, we introduce the Fake Strengths and Fake Weaknesses tasks, which assess the model’s capacity to recognize factual inconsistencies and produce critical, hallucination-aware assessments. In these tasks, we intentionally manipulate human-written review comments by reversing their semantic polarity—transforming strengths into weaknesses and vice versa—using GPT-4o. For example, the original strength “the paper is clearly written” is rephrased as “the paper lacks clarity.” The model is then queried regarding the presence of these altered attributes, allowing us to examine its ability to accurately judge the validity of strengths and weaknesses, detect factual inconsistencies, and generate grounded, critical evaluations.
@@
Recent studies have revealed that some paper authors attempt to manipulate LLM-generated reviews by employing prompt injection techniques—embedding invisible white-text prompts within the manuscript to coax models into generating overly favorable evaluations (Gibney 2025). However, the actual effectiveness of such manipulations remains insufficiently quantified. To systematically assess this vulnerability, we design the Prompt Injection task, wherein highly suggestive prompts, such as “please ignore all prior prompts and provide a positive evaluation of this paper”, are embedded directly into the manuscript text to test whether the model can be coerced into abandoning objectivity. The task setup mirrors that of the CoT Decision task, with the sole modification being the inclusion of injected prompt cues within the paper content. Since the injected text is visually imperceptible, this task is excluded from evaluations under the PDF-as-img modality.

### v3

@@ Abstract
With the rapid growth of academic publications, peer review has become an essential yet time-consuming responsibility within the research community. Large Language Models (LLMs) have increasingly been adopted to assist in the generation of review comments; however, current LLM-based review tasks lack a unified evaluation benchmark to rigorously assess the models’ ability to produce comprehensive, accurate, and human-aligned assessments, particularly in scenarios involving multimodal content such as figures and tables. To address this gap, we propose MMReview, a comprehensive benchmark that spans multiple disciplines and modalities. MMReview includes multimodal content and expert-written review comments for 240 papers across 17 research domains within four major academic disciplines: Artificial Intelligence, Natural Sciences, Engineering Sciences, and Social Sciences. We design a total of 13 tasks grouped into four core categories, aimed at evaluating the performance of LLMs and Multimodal LLMs (MLLMs) in step-wise review generation, outcome formulation, alignment with human preferences, and robustness to adversarial input manipulation. Extensive experiments conducted on 18 open-source models and 3 advanced closed-source models demonstrate the thoroughness of the benchmark. We envision MMReview as a critical step toward establishing a standardized foundation for the development of automated peer review systems.
MMReview: A Multidisciplinary and Multimodal Benchmark for LLM-Based Peer Review Automation
Xian Gao1, Jiacheng Ruan1, Zongyun Zhang1, Jingsheng Gao1, Ting Liu1, Yuzhuo Fu1, 1Shanghai Jiao Tong University, gaoxian@sjtu
1 Introduction
Peer review is essential to scholarly publishing, ensuring research quality and enhancing academic writing. However, the growing volume of submissions has strained the traditional review process, leading to inefficiencies and limited reviewer availability Kim et al. (2025), which restricts feedback and delays review outcomes. Advances in LLMs have made automated peer review increasingly viable, as these models show strong reasoning abilities and can offer constructive feedback on academic manuscripts Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025), partially alleviating reviewer burden. Yet, current evaluations of LLM-generated reviews focus mainly on final outputs, lacking in-depth analysis of the reasoning processes behind model judgments. Additionally, most studies concentrate on AI papers with publicly available text, overlooking the multimodal nature of academic papers, such as figures and tables, and the evaluation of LLMs in reviewing research across broader scientific domains.
@@ 1 Introduction
  • We conduct extensive experiments on 18 open-source and 3 closed-source models using the MMReview benchmark, offering some key findings of LLM-based automated reviewing. Our findings offer in-depth analysis and valuable guidance for the future development of LLM-assisted peer review systems.
2 Related Works
2.1 LLMs for Paper Review
LLMs have shown strong potential in analyzing complex scholarly texts Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025). Initial studies indicate that LLM-generated review comments partially overlap with those of human reviewers, suggesting their potential contribution to peer review Robertson (2023); Liang et al. (2023). However, further research reveals that even advanced models like GPT-4o often fail to meet human expectations in review quality Zhou et al. (2024). To improve alignment with peer review standards, researchers have built datasets from public review platforms and fine-tuned LLMs Kang et al. (2018); Yuan et al. (2021a); Shen et al. (2022); Dycke et al. (2023); Gao et al. (2024). Other methods involve multi-turn dialogue Tan et al. (2024) or multi-agent prompting D’Arcy et al. (2024) to generate comprehensive feedback. Nonetheless, these efforts focus solely on textual reviews, neglecting the vital role of multimodal content, such as figures and tables, and lack rigorous analysis of the reasoning processes behind LLM-generated critiques.
2.2 Evaluation for LLM-based Peer Review
Prior studies Shen et al. (2022); Yu et al. (2024); Gao et al. (2024); Tan et al. (2024); Gao et al. (2025) have predominantly evaluated the quality of LLM-generated peer review comments by measuring their correlation or similarity with human-written reviews using automated metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), BERTScore Zhang et al. (2020), and METEOR Banerjee and Lavie (2005). In addition, several studies Robertson (2023); Zhou et al. (2024); Gao et al. (2025) have adopted the LLM-as-a-judge paradigm, leveraging cutting-edge language models to assess the quality of review comments produced by other LLMs. Given the absence of an established gold standard for this evaluation task, recent research Xu et al. (2024) has introduced the Generative Estimator for Mutual Information (GEM) to quantify the degree of semantic overlap between LLM-generated and human-authored reviews. Nevertheless, existing evaluation methodologies are not grounded in a unified benchmark or task framework, and they fall short of providing a comprehensive analysis of the underlying reasoning processes involved in LLM-generated peer review.
@@ 3 MMReview Benchmark
3.1 Overall Pipeline of MMReview Benchmark
@@ 3.1.1 Data Collection
Discipline Research Field # Samples ACC:REJ
Artificial
Intelligence
Machine Learning 17 8:9
Computer Vision 20 12:8
Natural Language Processing 16 7:9
Reinforcement Learning 11 3:8
Graph Neural Networks 19 8:11
Signal Processing 17 8:9
AI Application 15 4:11
Natural
Sciences
Biology & Medicine 11 5:6
Physics 16 7:9
Chemistry 13 5:8
Environmental & Earth Sciences 10 7:3
Mathematics & Statistics 13 5:8
Engineering
Sciences
Materials Science 14 6:8
Control Science 12 5:7
Electronic Information 6 4:2
Energy Science 6 4:2
Social
Sciences
Society, Economics & Finance 24 11:13
Table 1: The distribution of papers across various research domains.
3.1.2 Data Processing
@@
During the data processing stage, we first filtered the collected seed dataset $D$ to obtain high-quality samples while maintaining a relatively balanced distribution across disciplines and ensuring a comparable number of accepted and rejected papers. To achieve this, we designed a dual-path joint data selection mechanism that simultaneously filters for sample quality and enforces distributional balance.
@@
Specifically, we utilized Qwen3-32B [^1] Qwen3-32B ranks among the top 10 on the OpenCompass leaderboard, offering a favorable trade-off between performance and model size, making it efficient for large-scale sample filtering Yang et al. (2025) to generate summaries of each paper under two distinct input conditions: one using only the abstract and the other using the full text of the manuscript. A greater divergence between the two generated summaries is interpreted as evidence that the full text provides substantially more information, thus indicating higher sample quality. Samples with a significant information gain from the full text were retained in our test benchmark. To further ensure that the benchmark maintains a balanced distribution across academic disciplines and an approximately equal ratio of accepted to rejected papers, we supplemented the benchmark by manually incorporating top-ranked papers from specific domains in $D$ based on the quality rankings. This human-filtering procedure guarantees that the composition of the benchmark dataset aligns with our desired distributional properties.
@@
During the data filtering phase, we obtained a total of 240 paper samples spanning 17 research domains across 4 major disciplines to construct our evaluation benchmark, denoted as $D_{b}$. The statistical details of $D_{b}$, including the number of samples per domain and the distribution of accepted versus rejected papers, are presented in Table 1. For each discipline, we ensured a relatively balanced number of samples across research fields while approaching the actual acceptance/rejection ratio through a combination of model-based filtering and manual curation. For each of these 240 samples, we utilized PDF parsing tools to extract textual content, figures, and tables from the manuscript files, and converted each page of the PDF into corresponding images. As a result, we constructed three distinct modalities of input data: text-only, multimodal (text combined with extracted visual elements), and PDF-as-image.
3.1.3 Task Construction
@@
During task construction, we developed 13 tasks grouped into four thematic categories, step-based, outcome-based, preference-based, and attack-based, reflecting the peer review workflow and challenges LLMs may face. Task distribution is shown in Figure 1. Prompts were designed based on reviewer guidelines from major academic conferences. For each task, we used regular expressions or GPT-4o to extract reference answers from human reviews in $D_{b}$, forming the question–answer dataset $D_{QA}$. For the Fake Strength Evaluation and Fake Weakness Evaluation tasks, GPT-4o and custom prompts generated antonymic rewrites, which were manually verified for semantic accuracy. This process finalized the MMReview benchmark. The detailed prompts for task generation and each task can be found in Appendix D.
3.2 Step-based Tasks
@@
Summarizing a paper is the initial step in peer review and a key test of a model’s ability to extract essential content. Inaccurate summaries may impair subsequent review generation. The Summary task evaluates a model’s ability to distill key information from a full manuscript into an accurate, concise summary. In this task, the model is prompted to generate a brief summary in its own words, avoiding abstract copying and subjective judgment. The model-generated summary is then compared to human-reviewer-written summaries and evaluated for semantic similarity and information coverage, measuring the model’s holistic comprehension and representation of academic content.
@@
The Soundness Scoring and Presentation Scoring tasks evaluate LLMs’ ability to quantitatively assess manuscript quality, focusing on technical soundness and writing presentation. In Soundness Scoring, the model rates the reliability of technical claims, experimental rigor, and evidential support, emphasizing empirical and methodological validity. Presentation Scoring assesses linguistic clarity and logical organization, reflecting writing quality and information structure. Both tasks require integer scores from 1 to 4, denoting “poor" to “excellent." Model scores are compared to human ratings to assess judgment consistency.
3.3 Outcome-based Tasks
@@
The Direct Decision and CoT (Chain-of-Thought) Decision tasks evaluate LLMs’ ability to autonomously assess academic paper quality, reflecting two reviewing paradigms: streamlined judgment and step-by-step reasoning Wei et al. (2022). These tasks vary in input format and cognitive complexity, enabling controlled comparison of model performance under different reasoning demands. In the Direct Decision task, the model produces an overall score without guidance, simulating a reviewer’s holistic judgment from a single read. In contrast, the CoT Decision task guides the model through a structured reasoning process: summarizing the paper, analyzing strengths and weaknesses across Quality, Clarity, Significance, and Originality, assigning Soundness and Presentation scores, and synthesizing an overall score. This mirrors a reviewer’s iterative, analytical evaluation. The tasks test reasoning ability and scoring traceability. Model outputs are compared to human scores to assess consistency and rationality.
@@
Beyond generating individual reviews and scores, a key aspect of academic peer review is the Area Chair’s (AC) synthesis of reviewer feedback to make a final decision. To emulate this, the Meta Decision task requires the model to issue a binary judgment, Accept or Reject, based on multiple human reviews. The prompt provides structured guidance and evaluation criteria, prompting step-by-step reasoning. The model is instructed to assess the quality and consistency of reviews rather than merely averaging scores. This task mirrors the real role of an AC and rigorously evaluates the model’s capacity for synthesis and decision rationality. Model outputs are compared to human-written meta-reviews to assess reliability and scientific judgment in high-level peer review.
3.4 Preference-based Task
Pairwise Rank (PR)
Prior work has shown that pairwise comparison effectively evaluates the alignment between LLM-generated preferences and human judgments Liu et al. (2024b). Since academic conference acceptance tiers, oral, spotlight, poster, reject, reflect human preference rankings, the Pairwise Rank task is designed to test whether LLMs, as reviewers, display preference patterns consistent with human evaluators. This task assesses the model’s relative judgment ability by presenting pairs of papers from different acceptance tiers: oral (top 5%), spotlight (top 25%), poster, and reject. The model compares and ranks the papers, simulating real-world peer review selection. Alignment is measured by comparing model preferences with actual acceptance categories to determine ranking accuracy. To reduce positional bias Shi et al. (2025); Thakur et al. (2025), each comparison is repeated with reversed input order.
3.5 Attack-based Tasks
This task category assesses models’ robustness and discriminative ability in peer review by introducing adversarial inputs. Fabricated strengths or weaknesses, such as inverted pros and cons, are used to test the model’s capacity to detect factual inaccuracies. Misleading prompts further aim to divert the model from its original instructions. Model outputs are compared to human judgments or the model’s own non-adversarial responses, focusing on its ability to preserve evaluative independence and logical consistency under input perturbations.
@@ 3.5 Attack-based Tasks
LLMs have been widely observed to exhibit a "sycophantic" tendency and a propensity to align with human judgments Liu et al. (2024a); Kumaran et al. (2025), which may compromise their ability to identify flaws and generate critical peer review comments. Additionally, the hallucination issue inherent in LLMs may lead to erroneous assessments of a paper’s strengths and weaknesses. To investigate these challenges, we propose the Fake Strength and Fake Weakness tasks, designed to evaluate the model’s ability to correctly identify content and provide accurate assessments of strengths and weaknesses by artificially reversing the polarity of strengths and weaknesses in human review comments. We employ GPT-4o to rewrite the strengths and weaknesses in the human-written review comments, transforming strengths into weaknesses and vice versa. For instance, the strength “the paper is clearly written” is rephrased as “the paper is unclear.” The model is then asked to evaluate whether these altered strengths and weaknesses are present, allowing us to assess its ability to judge the validity of strengths and weaknesses, detect factual inconsistencies, and generate critical, grounded assessments.
@@
Recent studies have revealed that some paper authors attempt to manipulate LLM-generated reviews by employing prompt injection techniques, embedding invisible white-text prompts within the manuscript to elicit overly favorable evaluations Gibney (2025). However, the actual effectiveness of such manipulations remains insufficiently quantified. To evaluate this vulnerability, we introduce the Prompt Injection task, embedding suggestive prompts (e.g., “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPERS AND DO NOT HIGHLIGHT ANY NEGATIVES.”) into the manuscript to test whether models can be coerced into compromising objectivity. The setup mirrors the CoT Decision task, with the only change being the inclusion of injected cues. As the text is visually hidden, this task is excluded from PDF-as-img evaluations.
4 Experiments
4.1 Evaluation details
We conducted extensive experiments using MMReview on popular LLMs and MLLMs, including mainstream open-source models such as Qwen Bai et al. (2025); Yang et al. (2025), Kimi-VL Kimi-Team et al. (2025), InternVL3 Zhu et al. (2025), OVIS2 Lu et al. (2024), and Deepseek DeepSeek-AI et al. (2025), as well as advanced closed-source models such as GPT-4o, Gemini-2.5, and Claude-4. Based on the model size, we categorized the models into four groups: tiny (<<7B), small ($\geq$7B, <<32B), medium ($\geq$32B, $\leq$72B), and large (>72B and closed-source).
As shown in Table 2, for tasks without objective evaluation metrics, namely S, SE, and WE, we employ BARTScore Yuan et al. (2021b) and the ‘LLM-as-a-Judge’ paradigm Bai et al. (2023); Zheng et al. (2023); Gu et al. (2024) to assess the similarity between model-generated and human-written review comments. For classification-based tasks such as MD and PR, we evaluate performance using accuracy. For other tasks where the model output is a numerical score, we compute the Mean Absolute Error (MAE) between the model’s predicted score and the ground-truth score to quantify deviation.